# mOSCAR: A Large-scale Multilingual and Multimodal Document-level Corpus

### v1

@@ Appendix A Appendix
Acehnese ace_Latn Austronesian Latin 7,803 32,461 2,889,134
Mesopotamian Arabic acm_Arab Afro-Asiatic Arabic 2,274 10,620 1,047,748
Tunisian Arabic aeb_Arab Afro-Asiatic Arabic 7,640 41,570 2,715,187
Afrikaans afr_Latn Indo-European Latin 54,895 247,774 39,956,585
South Levantine Arabic ajp_Arab Afro-Asiatic Arabic 12,098 87,837 5,167,813
Tosk Albanian als_Latn Indo-European Latin 861,678 2,569,164 452,737,251
Amharic amh_Ethi Afro-Asiatic Ge‘ez 39,588 152,646 35,089,019
North Levantine Arabic apc_Arab Afro-Asiatic Arabic 19,904 128,966 9,560,701
Modern Standard Arabic arb_Arab Afro-Asiatic Arabic 3,936,851 15,126,931 3,401,919,964
Najdi Arabic ars_Arab Afro-Asiatic Arabic 60,229 296,741 43,610,873
Moroccan Arabic ary_Arab Afro-Asiatic Arabic 142,386 698,051 204,723,454
Egyptian Arabic arz_Arab Afro-Asiatic Arabic 835,529 4,054,632 653,626,387
Assamese asm_Beng Indo-European Bengali 3,948 9,210 640,390
Asturian ast_Latn Indo-European Latin 165,745 962,723 37,547,944
Awadhi awa_Deva Indo-European Devanagari 29,324 107,483 4,961,635
Central Aymara ayr_Latn Aymaran Latin 27,384 151,889 5,148,970
South Azerbaijani azb_Arab Turkic Arabic 8,274 38,233 5,256,693
North Azerbaijani azj_Latn Turkic Latin 516,021 1,808,060 257,825,849
Bashkir bak_Cyrl Turkic Cyrillic 4,532 17,174 3,038,766
Bambara bam_Latn Manding Latin 7,674 39,190 1,243,332
Balinese ban_Latn Austronesian Latin 1,886 11,266 542,015
Belarusian bel_Cyrl Indo-European Cyrillic 63,309 287,539 72,976,520
Bemba bem_Latn Atlantic–Congo Latin 1,096 7,479 1,340,471
Bengali ben_Beng Indo-European Bengali 270,406 947,035 35,858,814
Bhojpuri bho_Deva Indo-European Devanagari 6,366 28,131 875,463
Banjar bjn_Latn Austronesian Latin 5,427 27,803 1,898,526
Bosnian bos_Latn Indo-European Latin 1,960,599 7,633,049 1,255,000,505
Buginese bug_Latn Austronesian Latin 3,312 18,648 588,678
Bulgarian bul_Cyrl Indo-European Cyrillic 2,591,998 11,670,028 1,760,971,620
Catalan cat_Latn Indo-European Latin 1,153,864 4,736,634 606,447,390
Cebuano ceb_Latn Austronesian Latin 16,990 91,234 10,748,818
Czech ces_Latn Indo-European Latin 3,918,837 13,291,309 2,823,172,996
Central Kurdish ckb_Arab Indo-European Arabic 36,725 136,566 22,322,689
Crimean Tatar crh_Latn Turkic Latin 6,376 24,124 1,742,727
Welsh cym_Latn Indo-European Latin 40,408 165,897 27,748,345
Danish dan_Latn Indo-European Latin 2,076,298 9,559,600 1,238,277,499
German deu_Latn Indo-European Latin 20,662,696 87,976,200 8,544,986,218
Southwestern Dinka dik_Latn Nilo-Saharan Latin 1,712 6,635 1,319,943
Greek ell_Grek Indo-European Greek 4,916,081 15,209,058 2,923,201,041
English eng_Latn Indo-European Latin 52,215,013 207,904,315 33,570,108,782
Esperanto epo_Latn Artificial Latin 25,157 124,996 28,586,195
Estonian est_Latn Uralic Latin 1,040,368 5,217,366 619,215,048
Basque eus_Latn Isolate Latin 849,043 3,445,539 277,145,498
Faroese fao_Latn Indo-European Latin 15,411 60,340 6,691,327
Fijian fij_Latn Austronesian Latin 1,528 8,776 487,388
Finnish fin_Latn Uralic Latin 2,396,033 10,365,333 1,781,044,864
French fra_Latn Indo-European Latin 20,305,739 78,179,601 14,362,579,829
Friulian fur_Latn Indo-European Latin 37,290 256,456 5,949,600
Nigerian Fulfulde fuv_Latn Atlantic-Congo Latin 1,568 7,124 401,852
West Central Oromo gaz_Latn Afro-Asiatic Latin 4,058 11,763 1,786,093
Scottish Gaelic gla_Latn Indo-European Latin 29,710 153,249 14,605,090
Irish gle_Latn Indo-European Latin 68,858 315,132 47,438,400
Galician glg_Latn Indo-European Latin 518,973 2,381,475 217,063,180
Guarani grn_Latn Tupian Latin 490,945 2,416,633 89,921,114
Gujarati guj_Gujr Indo-European Gujarati 23,062 91,320 3,324,866
Haitian Creole hat_Latn Indo-European Latin 257,745 1,570,699 62,847,106
Hausa hau_Latn Afro-Asiatic Latin 25,364 104,934 13,089,932
Hebrew heb_Hebr Afro-Asiatic Hebrew 1,109,591 4,766,483 893,327,320
Hindi hin_Deva Indo-European Devanagari 579,430 1,830,667 122,558,353
Chhattisgarhi hne_Deva Indo-European Devanagari 1,581 7,263 273,174
Croatian hrv_Latn Indo-European Latin 1,719,617 8,425,510 1,010,674,096
Hungarian hun_Latn Uralic Latin 3,534,506 15,390,083 2,831,715,050
Armenian hye_Armn Indo-European Armenian 339,962 1,141,885 205,635,952
Igbo ibo_Latn Atlantic-Congo Latin 11,529 68,049 8,701,070
Ilocano ilo_Latn Austronesian Latin 78,872 523,195 8,116,113
Indonesian ind_Latn Austronesian Latin 7,016,291 17,324,777 3,981,843,468
Icelandic isl_Latn Indo-European Latin 244,676 1,027,465 137,015,973
Italian ita_Latn Indo-European Latin 12,937,153 47,476,971 8,311,790,842
Javanese jav_Latn Austronesian Latin 24,785 135,583 16,908,805
Japanese jpn_Jpan Japonic Kanji 14,415,292 23,893,768 8,923,348,944
Kabyle kab_Latn Afro-Asiatic Latin 18,508 106,730 4,079,553
Kannada kan_Knda Dravidian Kannada 12,978 42,621 1,442,776
Kashmiri kas_Arab Indo-European Arabic 3,109 11,408 5,731,910
Georgian kat_Geor Kartvelian Georgian 354,436 1,304,281 275,223,026
Kazakh kaz_Cyrl Turkic Cyrillic 252,242 732,648 140,049,214
Halh Mongolian khk_Cyrl Mongolic Cyrillic 124,412 508,217 84,535,241
Khmer khm_Khmr Austroasiatic Kher 24,495 122,243 3,043,925
Kinyarwanda kin_Latn Atlantic-Congo Latin 30,401 172,201 12,049,616
Kyrgyz kir_Cyrl Uralic Cyrillic 53,010 199,713 34,404,281
Northern Kurdish kmr_Latn Indo-European Latin 39,262 164,666 23,834,960
Korean kor_Hang Koreanic Hanja 2,614,089 13,563,283 2,006,080,705
Lao lao_Laoo Kra-Dai Lao 50,611 208,768 31,029,380
Ligurian lij_Latn Indo-European Latin 8,751 56,266 2,958,179
Limburgish lim_Latn Indo-European Latin 189,547 1,076,047 42,534,327
Lingala lin_Latn Atlantic-Congo Latin 24,614 152,132 4,053,459
Lithuanian lit_Latn Indo-European Latin 1,688,811 8,869,443 1,161,476,040
Lombard lmo_Latn Indo-European Latin 30,506 151,855 9,058,614
Latgalian ltg_Latn Indo-European Latin 11,948 61,624 4,148,492
Luxembourgish ltz_Latn Indo-European Latin 44,987 246,346 16,676,872
Ganda lug_Latn Afro-Asiatic Latin 1,878 7,215 789,917
Mizo lus_Latn Sino-Tibetan Latin 7,880 26,817 4,978,472
Standard Latvian lvs_Latn Indo-European Latin 896,243 4,141,648 587,653,855
Magahi mag_Deva Indo-European Devanagari 1,097 3,847 205,763
Malayalam mal_Mlym Dravidian Malayalam 14,140 52,679 1,689,010
Marathi mar_Deva Indo-European Devanagari 50,391 163,868 6,689,250
Minangkabau min_Latn Austronesian Latin 9,341 35,309 1,256,931
Macedonian mkd_Cyrl Indo-European Cyrillic 542,250 1,853,070 307,232,151
Maltese mlt_Latn Afro-Asiatic Latin 120,888 709,242 36,097,957
Maori mri_Latn Austronesian Latin 24,322 130,137 24,957,914
Burmese mya_Mymr Sino-Tibetan Mon 8,144 44,188 539,527
Dutch nld_Latn Indo-European Latin 17,096,727 65,606,013 9,670,041,731
Norwegian Nynorsk nno_Latn Indo-European Latin 199,355 1,012,313 67,799,774
Norwegian Bokmål nob_Latn Indo-European Latin 2,229,702 9,698,128 1,294,178,095
Nepali npi_Deva Indo-European Devanagari 31,239 127,193 3,138,539
Nyanja nya_Latn Atlantic-Congo Latin 12,047 67,192 8,596,769
Occitan oci_Latn Indo-European Latin 164,852 671,881 59,309,549
Odia ory_Orya Indo-European Odia 4,319 15,574 378,635
Pangasinan pag_Latn Austronesian Latin 4,214 32,287 546,071
Eastern Panjabi pan_Guru Indo-European Gurmukhi 11,497 46,168 1,887,991
Papiamento pap_Latn Indo-European Latin 55,224 363,015 10,002,655
Southern Pasto pbt_Arab Indo-European Arabic 32,604 110,807 29,170,322
Western Persian pes_Arab Indo-European Arabic 7,048,946 25,200,571 6,210,479,015
Plateau Malgasy plt_Latn Austronesian Latin 32,521 120,673 29,263,848
Polish pol_Latn Indo-European Latin 14,549,605 60,639,244 11,104,144,109
Portuguese por_Latn Indo-European Latin 8,145,664 26,530,423 4,760,063,083
Dari prs_Arab Indo-European Arabic 515,041 2,589,859 517,053,967
Ayacucho Quechua quy_Latn Quechuan Latin 1,578 11,817 362,690
Romanian ron_Latn Indo-European Latin 5,180,171 17,964,048 3,548,291,261
Rundi run_Latn Atlantic-Congo Latin 20,001 67,096 8,686,054
Russian rus_Cyrl Indo-European Cyrillic 15,913,845 69,542,828 18,909,213,208
Sango sag_Latn Atlantic-Congo Latin 2,124 13,556 454,455
Sicilian scn_Latn Indo-European Latin 73,199 424,362 27,110,743
Sinhala sin_Sinh Indo-European Sinhalese 58,767 221,183 14,270,972
Slovak slk_Latn Indo-European Latin 3,008,599 15,067,234 1,963,804,563
Slovenian slv_Latn Indo-European Latin 1,472,025 7,210,285 935,834,754
Samoan smo_Latn Austronesian Latin 12,346 71,359 14,954,824
Shona sna_Latn Atlantic-Congo Latin 12,698 68,782 6,112,600
Sindhi snd_Arab Indo-European Arabic 21,095 74,289 17,647,825
Somali som_Latn Afro-Asiatic Latin 77,343 301,429 34,554,975
Southern Sotho sot_Latn Atlantic-Congo Latin 7,718 43,146 6,156,450
Spanish spa_Latn Indo-European Latin 22,713,366 78,361,087 14,616,773,475
Sardinian srd_Latn Indo-European Latin 675,539 4,059,493 106,159,957
Serbian srp_Cyrl Indo-European Cyrillic 604,557 2,286,171 401,223,741
Sundanese sun_Latn Austronesian Latin 44,310 236,025 13,627,832
Swedish swe_Latn Indo-European Latin 3,302,730 10,860,518 1,779,284,152
Swahili swh_Latn Atlantic-Congo Latin 137,134 593,418 59,454,896
Silesian szl_Latn Indo-European Latin 23,535 132,459 5,996,972
Tamil tam_Taml Dravidian Tamil 36,196 167,669 4,834,946
Tatar tat_Cyrl Turkic Cyrillic 37,188 143,842 22,831,350
Telugu tel_Telu Dravidian Telugu 22,974 81,033 2,273,772
Tajik tgk_Cyrl Turkic Cyrillic 125,236 417,591 90,503,778
Tagalog tgl_Latn Austronesian Latin 151,437 673,814 97,708,639
Thai tha_Thai Kra-Dai Thai 2,983,837 11,621,786 2,839,211,104
Tigrinya tir_Ethi Afro-Asiatic Ge‘ez 2,657 8,707 1,725,422
Tok Pisin tpi_Latn Indo-European Latin 5,063 35,169 460,853
Turkmen tuk_Latn Turkic Latin 13,024 57,354 9,766,999
Turkish tur_Latn Turkic Latin 4,478,700 12,401,091 2,394,669,068
Twi twi_Latn Atlantic-Congo Latin 3,305 13,634 495,220
Uyghur uig_Arab Turkic Arabic 10,713 41,709 6,785,318
Ukrainian ukr_Cyrl Indo-European Cyrillic 2,721,424 10,929,796 1,928,351,595
Urdu urd_Arab Indo-European Arabic 407,098 1,239,125 242,007,283
Northern Uzbek uzn_Latn Turkic Latin 156,632 798,155 89,022,562
Venetian vec_Latn Indo-European Latin 330,611 1,830,777 71,077,531
Vietnamese vie_Latn Viet-Muong Latin 12,621,521 47,411,488 11,616,191,199
Wolof wol_Latn Atlantic-Congo Latin 4,658 20,380 1,596,432
Xhosa xho_Latn Atlantic-Congo Latin 25,950 142,387 15,809,823
Eastern Yiddish ydd_Hebr Indo-European Hebrew 12,486 57,510 17,369,727
Yoruba yor_Latn Atlantic-Congo Latin 56,700 286,933 32,614,558
Yue Chinese yue_Hant Sino-Tibetan Hant 33,671 203,513 24,172,441
Chinese (Simplified) zho_Hans Sino-Tibetan Hanzi 9,861,262 36,152,754 8,078,842,701
Chinese (Traditional) zho_Hant Sino-Tibetan Hant 3,967,966 16,307,258 2,962,854,441
Standard Malay zsm_Latn Austronesian Latin 1,179,744 5,488,632 432,667,199
Zulu zul_Latn Atlantic-Congo Latin 30,717 156,639 11,345,288

### v2

@@ A.3 Examples of documents
Acehnese ace_Latn Austronesian Latin 2,159 9,026 1,395,381
Mesopotamian Arabic acm_Arab Afro-Asiatic Arabic 1,282 5,621 704,549
Tunisian Arabic aeb_Arab Afro-Asiatic Arabic 5,933 34,270 2,308,455
Afrikaans afr_Latn Indo-European Latin 50,061 211,876 38,761,504
South Levantine Arabic ajp_Arab Afro-Asiatic Arabic 8,603 69,051 3,869,688
Tosk Albanian als_Latn Indo-European Latin 856,144 2,543,758 441,244,377
Amharic amh_Ethi Afro-Asiatic Ge‘ez 39,031 149,739 33,768,732
North Levantine Arabic apc_Arab Afro-Asiatic Arabic 16,198 110,792 8,268,237
Modern Standard Arabic arb_Arab Afro-Asiatic Arabic 3,794,792 14,757,353 3,346,786,610
Najdi Arabic ars_Arab Afro-Asiatic Arabic 52,102 261,275 39,066,487
Moroccan Arabic ary_Arab Afro-Asiatic Arabic 117,957 584,301 188,462,338
Egyptian Arabic arz_Arab Afro-Asiatic Arabic 761,113 3,785,164 635,018,784
Assamese asm_Beng Indo-European Bengali 2,947 7,228 543,676
Asturian ast_Latn Indo-European Latin 87,649 533,723 25,499,269
Awadhi awa_Deva Indo-European Devanagari 8,179 29,142 2,293,620
Central Aymara ayr_Latn Aymaran Latin 10,112 57,294 2,343,403
South Azerbaijani azb_Arab Turkic Arabic 3,411 14,825 3,143,946
North Azerbaijani azj_Latn Turkic Latin 511,832 1,796,046 256,160,442
Bashkir bak_Cyrl Turkic Cyrillic 3,287 12,031 2,600,135
Bambara bam_Latn Manding Latin 3,011 17,666 446,961
Balinese ban_Latn Austronesian Latin 787 4,894 392,978
Belarusian bel_Cyrl Indo-European Cyrillic 60,443 276,672 71,854,171
Bemba bem_Latn Atlantic–Congo Latin 582 3,018 1,021,026
Bengali ben_Beng Indo-European Bengali 204,475 758,222 30,400,395
Bhojpuri bho_Deva Indo-European Devanagari 4,190 18,339 715,786
Banjar bjn_Latn Austronesian Latin 1,764 9,017 1,093,443
Bosnian bos_Latn Indo-European Latin 635,750 2,642,491 423,073,661
Buginese bug_Latn Austronesian Latin 584 2,379 167,459
Bulgarian bul_Cyrl Indo-European Cyrillic 2,578,191 11,601,214 1,736,106,287
Catalan cat_Latn Indo-European Latin 1,132,056 4,638,966 598,942,711
Cebuano ceb_Latn Austronesian Latin 14,924 75,258 10,221,371
Czech ces_Latn Indo-European Latin 3,736,126 12,683,461 2,767,295,966
Central Kurdish ckb_Arab Indo-European Arabic 36,413 135,461 21,622,335
Crimean Tatar crh_Latn Turkic Latin 2,744 10,079 1,173,321
Welsh cym_Latn Indo-European Latin 38,616 155,591 27,237,252
Danish dan_Latn Indo-European Latin 2,020,516 9,214,031 1,207,829,704
German deu_Latn Indo-European Latin 20,265,504 86,393,702 8,315,212,019
Southwestern Dinka dik_Latn Nilo-Saharan Latin 1,233 4,766 1,098,795
Greek ell_Grek Indo-European Greek 4,895,433 15,147,284 2,909,427,055
English eng_Latn Indo-European Latin 51,658,029 205,363,181 32,599,001,993
Esperanto epo_Latn Artificial Latin 23,619 112,577 26,976,847
Estonian est_Latn Uralic Latin 1,022,368 5,108,102 589,045,973
Basque eus_Latn Isolate Latin 682,599 2,914,120 259,930,954
Faroese fao_Latn Indo-European Latin 14,921 56,934 6,579,921
Fijian fij_Latn Austronesian Latin 1,039 4,039 416,670
Finnish fin_Latn Uralic Latin 2,377,155 10,263,171 1,749,904,041
French fra_Latn Indo-European Latin 19,963,542 76,851,982 13,818,099,493
Friulian fur_Latn Indo-European Latin 15,823 120,878 2,550,209
Nigerian Fulfulde fuv_Latn Atlantic-Congo Latin 919 4,281 264,234
West Central Oromo gaz_Latn Afro-Asiatic Latin 3,399 9,071 1,640,693
Scottish Gaelic gla_Latn Indo-European Latin 19,638 105,937 13,119,348
Irish gle_Latn Indo-European Latin 60,303 267,562 45,341,371
Galician glg_Latn Indo-European Latin 410,489 1,696,763 197,685,077
Guarani grn_Latn Tupian Latin 207,800 1,038,296 48,610,979
Gujarati guj_Gujr Indo-European Gujarati 21,916 87,805 3,202,096
Haitian Creole hat_Latn Indo-European Latin 105,777 667,801 34,261,838
Hausa hau_Latn Afro-Asiatic Latin 21,850 81,141 11,807,898
Hebrew heb_Hebr Afro-Asiatic Hebrew 1,098,800 4,708,947 859,238,720
Hindi hin_Deva Indo-European Devanagari 543,928 1,745,222 118,903,998
Chhattisgarhi hne_Deva Indo-European Devanagari 832 3,908 205,345
Croatian hrv_Latn Indo-European Latin 1,689,553 8,315,237 998,928,993
Hungarian hun_Latn Uralic Latin 3,515,058 15,293,132 2,811,446,583
Armenian hye_Armn Indo-European Armenian 336,285 1,126,920 199,883,484
Igbo ibo_Latn Atlantic-Congo Latin 7,089 41,672 3,014,602
Ilocano ilo_Latn Austronesian Latin 7,076 59,327 832,454
Indonesian ind_Latn Austronesian Latin 6,644,918 16,237,247 2,895,956,979
Icelandic isl_Latn Indo-European Latin 239,195 1,003,522 131,308,802
Italian ita_Latn Indo-European Latin 12,812,932 47,011,085 8,144,757,759
Javanese jav_Latn Austronesian Latin 18,192 100,952 15,206,708
Japanese jpn_Jpan Japonic Kanji 14,154,575 23,435,549 8,539,956,266
Kabyle kab_Latn Afro-Asiatic Latin 6,101 33,923 1,781,992
Kannada kan_Knda Dravidian Kannada 9,373 33,147 1,206,651
Kashmiri kas_Arab Indo-European Arabic 1,498 5,284 3,384,394
Georgian kat_Geor Kartvelian Georgian 353,471 1,300,710 274,042,522
Kazakh kaz_Cyrl Turkic Cyrillic 248,403 718,126 138,597,176
Halh Mongolian khk_Cyrl Mongolic Cyrillic 123,789 505,098 83,628,495
Khmer khm_Khmr Austroasiatic Kher 23,348 116,437 2,915,205
Kinyarwanda kin_Latn Atlantic-Congo Latin 20,381 108,280 10,268,334
Kyrgyz kir_Cyrl Uralic Cyrillic 51,221 194,092 33,981,180
Northern Kurdish kmr_Latn Indo-European Latin 34,593 142,634 21,972,155
Korean kor_Hang Koreanic Hanja 2,614,038 13,562,957 2,000,344,511
Lao lao_Laoo Kra-Dai Lao 49,925 205,452 30,098,274
Ligurian lij_Latn Indo-European Latin 3,581 26,740 1,046,463
Limburgish lim_Latn Indo-European Latin 70,099 443,903 25,465,590
Lingala lin_Latn Atlantic-Congo Latin 6,304 41,400 1,580,536
Lithuanian lit_Latn Indo-European Latin 1,673,790 8,772,570 1,153,604,941
Lombard lmo_Latn Indo-European Latin 14,053 61,359 6,270,646
Latgalian ltg_Latn Indo-European Latin 5,174 21,062 2,903,043
Luxembourgish ltz_Latn Indo-European Latin 27,946 142,470 13,925,521
Ganda lug_Latn Afro-Asiatic Latin 1,475 4,118 688,308
Mizo lus_Latn Sino-Tibetan Latin 7,009 22,630 4,106,536
Standard Latvian lvs_Latn Indo-European Latin 857,757 3,937,940 578,441,751
Magahi mag_Deva Indo-European Devanagari 290 1,088 94,031
Malayalam mal_Mlym Dravidian Malayalam 11,203 44,417 1,420,906
Marathi mar_Deva Indo-European Devanagari 43,720 142,001 6,164,176
Minangkabau min_Latn Austronesian Latin 1,523 7,300 447,320
Macedonian mkd_Cyrl Indo-European Cyrillic 539,149 1,841,846 304,592,615
Maltese mlt_Latn Afro-Asiatic Latin 56,666 327,331 27,114,870
Maori mri_Latn Austronesian Latin 20,840 114,680 24,524,962
Burmese mya_Mymr Sino-Tibetan Mon 6,575 36,661 406,016
Dutch nld_Latn Indo-European Latin 16,890,074 64,609,055 9,493,533,101
Norwegian Nynorsk nno_Latn Indo-European Latin 138,384 701,972 57,812,652
Norwegian Bokmål nob_Latn Indo-European Latin 2,192,012 9,534,178 1,267,421,216
Nepali npi_Deva Indo-European Devanagari 28,042 116,363 2,892,865
Nyanja nya_Latn Atlantic-Congo Latin 11,749 65,324 8,513,823
Occitan oci_Latn Indo-European Latin 61,681 323,632 21,029,975
Odia ory_Orya Indo-European Odia 3,759 14,373 340,695
Pangasinan pag_Latn Austronesian Latin 1,045 7,770 270,363
Eastern Panjabi pan_Guru Indo-European Gurmukhi 10,857 44,440 1,821,511
Papiamento pap_Latn Indo-European Latin 29,564 177,229 7,396,392
Southern Pashto pbt_Arab Indo-European Arabic 31,854 107,563 27,623,486
Western Persian pes_Arab Indo-European Arabic 6,995,368 24,998,370 6,061,794,870
Plateau Malgasy plt_Latn Austronesian Latin 32,119 119,506 28,542,084
Polish pol_Latn Indo-European Latin 14,492,239 60,362,860 10,994,239,010
Portuguese por_Latn Indo-European Latin 8,033,406 26,058,040 4,639,089,792
Dari prs_Arab Indo-European Arabic 421,097 2,101,038 399,037,437
Ayacucho Quechua quy_Latn Quechuan Latin 1,248 10,038 322,112
Romanian ron_Latn Indo-European Latin 5,131,444 17,790,793 3,484,865,185
Rundi run_Latn Atlantic-Congo Latin 17,798 55,060 8,140,230
Russian rus_Cyrl Indo-European Cyrillic 15,753,144 68,786,134 18,196,141,357
Sango sag_Latn Atlantic-Congo Latin 724 4,564 181,876
Sicilian scn_Latn Indo-European Latin 27,388 164,772 17,535,500
Sinhala sin_Sinh Indo-European Sinhalese 44,963 179,082 11,413,044
Slovak slk_Latn Indo-European Latin 2,979,681 14,894,160 1,951,406,321
Slovenian slv_Latn Indo-European Latin 1,456,026 7,106,291 928,101,642
Samoan smo_Latn Austronesian Latin 11,024 62,358 11,672,900
Shona sna_Latn Atlantic-Congo Latin 7,400 41,385 5,276,139
Sindhi snd_Arab Indo-European Arabic 20,615 70,992 16,686,668
Somali som_Latn Afro-Asiatic Latin 58,151 209,905 31,093,227
Southern Sotho sot_Latn Atlantic-Congo Latin 7,474 41,714 5,876,842
Spanish spa_Latn Indo-European Latin 22,218,630 76,372,709 13,882,047,139
Sardinian srd_Latn Indo-European Latin 336,476 2,220,976 68,281,992
Serbian srp_Cyrl Indo-European Cyrillic 593,332 2,251,042 394,477,097
Sundanese sun_Latn Austronesian Latin 16,438 89,379 9,549,957
Swedish swe_Latn Indo-European Latin 3,231,753 10,558,719 1,748,495,813
Swahili swh_Latn Atlantic-Congo Latin 96,770 365,792 52,827,863
Silesian szl_Latn Indo-European Latin 7,846 47,313 3,022,502
Tamil tam_Taml Dravidian Tamil 30,202 149,837 4,234,345
Tatar tat_Cyrl Turkic Cyrillic 34,489 133,014 22,255,423
Telugu tel_Telu Dravidian Telugu 16,107 54,100 1,633,579
Tajik tgk_Cyrl Turkic Cyrillic 119,383 395,470 87,519,228
Tagalog tgl_Latn Austronesian Latin 140,922 628,210 95,285,900
Thai tha_Thai Kra-Dai Thai 1,799,735 6,603,060 807,374,946
Tigrinya tir_Ethi Afro-Asiatic Ge‘ez 2,622 8,601 1,699,272
Tok Pisin tpi_Latn Indo-European Latin 785 5,888 97,298
Turkmen tuk_Latn Turkic Latin 12,372 54,002 9,650,172
Turkish tur_Latn Turkic Latin 4,448,111 12,304,912 2,356,627,784
Twi twi_Latn Atlantic-Congo Latin 286 2,041 78,227
Uyghur uig_Arab Turkic Arabic 10,614 41,367 6,602,690
Ukrainian ukr_Cyrl Indo-European Cyrillic 2,689,369 10,842,572 1,909,330,669
Urdu urd_Arab Indo-European Arabic 403,245 1,224,175 236,356,788
Northern Uzbek uzn_Latn Turkic Latin 113,772 581,861 81,808,833
Venetian vec_Latn Indo-European Latin 122,390 763,029 24,081,966
Vietnamese vie_Latn Viet-Muong Latin 12,296,989 46,339,341 11,462,111,787
Wolof wol_Latn Atlantic-Congo Latin 2,152 9,351 367,848
Xhosa xho_Latn Atlantic-Congo Latin 13,620 80,748 14,566,904
Eastern Yiddish ydd_Hebr Indo-European Hebrew 12,275 56,421 17,078,751
Yoruba yor_Latn Atlantic-Congo Latin 10,148 49,474 8,346,193
Yue Chinese yue_Hant Sino-Tibetan Hant 28,478 172,592 21,579,579
Chinese (Simplified) zho_Hans Sino-Tibetan Hanzi 8,326,440 29,575,591 5,199,137,981
Chinese (Traditional) zho_Hant Sino-Tibetan Hant 3,796,336 15,514,804 2,617,463,485
Standard Malay zsm_Latn Austronesian Latin 864,831 3,651,754 384,708,004
Zulu zul_Latn Atlantic-Congo Latin 13,089 73,167 9,654,461